# Rethinking Reuse in Dependency Supply Chains: Initial Analysis of NPM packages at the End of the Chain

### v2

@@ 4. Research Agenda
Our early results reveal five cases of packages that do not follow the majority mindset of ‘reuse as much as possible’. As a side effect, these packages may in-turn be more resilient to package vulnerabilities and seem outdated, but are still heavily depended upon. We split our research agenda into five dimensions to further investigate how to utilize how these packages reuse dependencies.
4.1. Characterizing End-of-Chain packages: A New Lens for Reuse
The characterization of end-of-chain packages is crucial to understanding their role in the software ecosystem and the strategic decisions around their reuse. These packages, with minimal dependencies, provide unique insights into how reuse can be optimized for stability and resilience in software projects.
@@ 4.1. Characterizing End-of-Chain packages: A New Lens for Reuse
    How can we characterize end-of-chain packages to better inform reuse decisions? Approach: Explore the unique characteristics of end-of-chain packages, such as their autonomy, low dependency footprint, and stability. Track the history to explore whether these packages once had dependencies that they eventually absorbed or removed.
@@
    What impact does removing or replacing end-of-chain packages have on the broader software ecosystem, and how should reuse strategies account for this? Approach: Investigate how the removal or replacement of end-of-chain packages affects projects that depend on them. We can study how upstream packages can be affected and whether the functionality can be easily absorbed by these upstream clients.
4.2. Strategic Dependency Management: When to Reuse or Not
To rethink reuse, it is essential to examine when and why dependencies should be reduced, absorbed, or replaced. Understanding the right timing and context for reusing packages or creating new solutions will optimize software performance, minimize risks, and improve long-term maintainability.
@@ 4.2. Strategic Dependency Management: When to Reuse or Not
    How can we optimize the decision-making process for when to reuse a package versus creating a new function? Approach: Research the trade-offs between using established packages and creating custom functions within the package.
@@
    What criteria should guide the decision to absorb (or bundle) dependencies into a package, and how can this inform the reuse strategy? Approach: Study the motivations behind bundling dependencies into a package (as seen with chalk) and how this affects the software ecosystem. Develop decision criteria that developers can use to decide whether to absorb or remove dependencies, taking into account long-term reuse and minimal dependency management.
4.3. Reimaging Supply Chain Resilience and Reuse Sustainability
The long-term sustainability and resilience of the software ecosystem depend on how packages are reused and maintained. Redefining reuse strategies that prioritize resilience and minimize risk can enhance the overall health of the ecosystem, especially in critical supply chains.
@@ 4.3. Reimaging Supply Chain Resilience and Reuse Sustainability
    How can we redefine reuse as a strategy for building resilience and sustainability in the software supply chain? Approach: Redefine reuse not just as a cost-saving strategy, but as a means of enhancing the resilience and sustainability of software ecosystems. Examine how minimal dependency and end-of-chain packages can play a more central role in resilient supply chains, and propose guidelines for their reuse in critical systems.
@@
    What factors contribute to the long-term sustainability of packages at the end of the chain, and how can this inform reuse policies? Approach: Study the factors that enable the long-term sustainability of end-of-chain packages, such as their lack of dependencies, community support, and focus on stability. Develop best practices for maintaining and reusing such packages over extended periods, considering both their technical and community aspects.
@@
    What are the trade-offs between reuse and ecosystem health, and how can we ensure that the reuse of packages does not negatively impact the ecosystem? Approach: Examine the ecological effects of widespread package reuse, including the risks of over-reliance on certain packages and the potential for cascading failures. Propose a balanced approach to reuse that ensures packages are used in ways that contribute to ecosystem health without causing dependency bloat or security vulnerabilities.
4.4. Redefining the Future of Reuse in a Minimal Dependency Ecosystem
As end-of-chain packages become more prominent, it is crucial to rethink how reuse is performed in an ecosystem with minimal dependencies. This redefinition can lead to more efficient, secure, and sustainable software architectures in the long run.
@@ 4.4. Redefining the Future of Reuse in a Minimal Dependency Ecosystem
    Can we develop a new model for package reuse that prioritizes self-sufficiency and reduces upstream dependency risk? Approach: Propose a new model for package reuse that encourages self-sufficiency, reduces risk exposure from upstream vulnerabilities, and focuses on reducing unnecessary dependencies. Explore how end-of-chain packages can serve as examples for creating more autonomous software components in the ecosystem.
@@
    What role do end-of-chain packages play in shaping the reuse culture within open-source communities, and how can this inform future practices? Approach: Explore how end-of-chain packages influence reuse culture within open-source communities. Investigate whether they promote a more sustainable, self-sufficient model of software development, and provide recommendations for fostering a reuse culture that values long-term stability and security over short-term convenience.
4.5. On Rethinking Reuse for the Ecosystem
Further investigation into the nuances of reuse strategies is essential to fully understand how minimal dependency ecosystems can thrive. The following questions focus on measuring impact, governance, and creating feedback mechanisms that foster sustainable reuse.
@@ 4.5. On Rethinking Reuse for the Ecosystem
    How can we measure the impact of reduced dependency on both the quality and performance of reused packages? Approach: Develop performance and quality metrics for evaluating packages that reduce their dependencies. Study how reducing dependencies can improve performance, security, and maintainability, and create guidelines for balancing these metrics when deciding to reuse a package.
@@
  3. (12)
    How can we create a feedback loop where the successful use of end-of-chain packages leads to broader adoption of minimal dependency strategies? Approach: Examine case studies of projects that successfully adopted end-of-chain packages and explore how their success can be used as a model for broader adoption of minimal dependency strategies. Develop a framework for creating feedback loops that encourage the wider adoption of these practices.
5. Discussion and Conclusion
Our results suggest that end-of-chain packages potentially play an important role in the NPM ecosystem, with four of the ten most depended upon packages on the dependency chain having no dependencies themselves. While there are obvious benefits to a lack of dependence on other software, such as a lack of inherited vulnerabilities, other benefits remain unknown. Additionally, we don’t know how these gains stack up against the productivity benefits of code reuse. By looking at five example cases of packages with no dependencies at the end of the supply chain, we make the case that there may be new insights to be gained from similar packages across software ecosystems. End-of-chain packages may have their own risk and rewards that need to considered, and the motivations behind the use of no dependencies are currently unstudied. We present a research agenda for understanding these aspects of end-of-chain packages.

### v3

@@ 5. Research Agenda
Our initial analysis reveals five cases of packages that do not follow the majority mindset of ‘reuse as much as possible’, but bring forth interesting questions about each library’s position in the software ecosystem. Informed from these cases, we split the research agenda into four dimensions to further investigate the case of end-of-chain packages.
5.1. Characterising End-of-Chain Packages
The characterization of end-of-chain packages is crucial to understanding their role in the software ecosystem and the strategic decisions around their reuse. These packages, with minimal dependencies, provide unique insights into how reuse can be optimized for stability and resilience in software projects. The aim is to explore the behavioural and structural characteristics of end-of-chain packages, such as size, update frequency and popularity.
@@ 5.1. Characterising End-of-Chain Packages
    Are core, highly-used packages more likely to have no dependencies? As we discovered with Case 1, some of the most popular packages in the NPM ecosystem have no dependencies. We aim to assess whether there is a statistical correlation between popularity measures such as download count, dependent count, etc. and having no dependencies.
@@
    What types of functionality are often implemented by end-of-chain packages? Many of the cases we investigated involved trivial, single purpose packages. These simple, limited scope libraries may lend themselves to having less dependencies. We ask whether certain problem domains are particularly likely to result in dependency-free implementations, either due to simplicity or to reduce attack surface.
@@
    Do end-of-chain libraries always begin with no dependencies, or do they emerge from reduction efforts? Some packages may have once used dependencies that were later removed. We seek to track their history to understand how and why dependency reduction occurs over time.
@@
    What techniques do developers use to reduce dependencies, and how prevalent are they (e.g. bundling, reimplementation)? We aim to understand the common techniques for reducing dependencies, based on the histories of existing end-of-chain packages.
5.2. Understanding Developer Motivations
In this section, we aim to delve deeper into the motivations of developers that both maintain or adopt end-of-chain libraries. The decision to eliminate dependencies may be influenced by concerns over security, package size, performance, or long-term maintainability. By understanding these choices, we can better frame the cultural, technical, and risk-based reasoning behind reuse decisions and ultimately provide more informed guidance for sustainable package development and adoption.
@@ 5.2. Understanding Developer Motivations
    What factors motivate developers to avoid dependencies? Developers may avoid dependencies to reduce security risks, minimise bloat, simplify maintenance, or retain control over code. We aim to investigate which motivations are most common and how they vary across contexts.
@@
    When and why do developers choose to bundle dependencies? In Case 5 we observe an example of bundling, with cited motivations such as speed. However, it may also obscure origin and trust boundaries. We aim to investigate when and why developers make these trade-offs.
@@
    To what extent do developers consider a package’s dependency chain when selecting what package to adopt? Are adoption decisions influenced by the presence or absence of dependencies, and if so, under what conditions? We explore whether dependency count plays a measurable role in decision-making for package adopters.
5.3. Evaluating the Risks, Impact and Benefits of No Dependencies
Removing dependencies can have risks and benefits. This section investigates the ecosystem-level implications of depending on—or eliminating—end-of-chain packages.
@@ 5.3. Evaluating the Risks, Impact and Benefits of No Dependencies
    What impact does removing or replacing end-of-chain packages have on the broader software ecosystem, and how should reuse strategies account for this? We aim to investigate how the removal or replacement of end-of-chain packages affects projects that depend on them. Looking at the example in Case 3, we can observe that some end-of-chain libraires may be candidates for removal. We can study how upstream packages can be affected and whether the functionality can be easily absorbed.
@@
    Do end-of-chain packages tend to require less long-term maintenance? Without dependency churn, these packages may require fewer updates (as in Case 2). We examine whether this correlates with reduced maintenance cost over time.
@@
    Are end-of-chain packages more secure? Having no dependencies may reduce the attack surface, but this alone may not guarantee safety. We investigate whether these packages demonstrate lower vulnerability rates in practice.
  4. (11)
    What are the risks and benefits of bundling dependencies, and how can this inform reuse strategy? Bundling may improve resilience and performance, but also reduce transparency. We examine its tradeoffs in real-world ecosystems. We aim to develop decision criteria that developers can use to decide whether to absorb or remove dependencies, taking into account long-term reuse and minimal dependency management.
  5. (12)
    When should developers reuse a package versus reduce dependency? Based on our taxonomy, we aim to identify cases where reuse adds clear value versus when it adds unnecessary risk or overhead, in order to optimize the decision makjing process for developers.
5.4. Rethinking Reuse for the Ecosystem
To rethink reuse, it is essential to examine the role of end-of-chain libraries, when and why dependencies should be reduced, and the risks and benefits. Informed by the previous research questions, we propose a redefining of reuse strategies that that prioritize resilience and minimize risk, in order to enhance the overall health of the ecosystem. This redefinition can lead to more efficient, secure, and sustainable software architectures in the long run. The following questions focus on measuring impact, governance, and creating feedback mechanisms that foster sustainable reuse.
  1. (13)
    Can we develop a new model for package reuse that prioritizes self-sufficiency and reduces upstream dependency risk? We aim to propose a new model for package reuse that encourages self-sufficiency, reduces risk exposure from upstream vulnerabilities, and focuses on reducing unnecessary dependencies. We will explore how end-of-chain packages can serve as examples for creating more autonomous software components in the ecosystem.
  2. (14)
    What factors contribute to the long-term sustainability of packages at the end of the chain, and how can this inform reuse policies? Study the factors that enable the long-term sustainability of end-of-chain packages, such as their lack of dependencies, community support, and focus on stability. Develop best practices for maintaining and reusing such packages over extended periods, considering both their technical and community aspects.
  3. (15)
    How can we measure the impact of reduced dependency on both the quality and performance of reused packages? Develop performance and quality metrics for evaluating packages that reduce their dependencies. Study how reducing dependencies can improve performance, security, and maintainability, and create guidelines for balancing these metrics when deciding to reuse a package.
@@ 5.4. Rethinking Reuse for the Ecosystem
6. Conclusion
In this paper, we identify the prevalence of end-of-chain packages within the dependency chain of the most commonly used NPM libraries, making up over half of dependencies in our sample. Through five motivating cases, we show that these packages are not homogeneous: some are core ecosystem infrastructure with active maintenance, some remain stable due to their simplicity, while others raise questions about triviality, risk, and sustainability. Together, these examples illustrate that dependency avoidance is not merely an accident of design, but can also be a conscious strategy.
We argue that understanding end-of-chain packages is essential to rethinking software reuse. Whereas the prevailing philosophy of “reuse as much as possible” has accelerated development, it has also introduced fragility into software supply chains. End-of-chain packages challenge this narrative by suggesting that in certain contexts, minimizing or eliminating dependencies can lead to greater sustainability and control. At the same time, these packages present their own trade-offs and potential risks, which remain understudied.
Looking forward, we present a clear research agenda: characterizing the traits of end-of-chain libraries across ecosystems, uncovering developer motivations for reducing dependencies, and evaluating the broader trade-offs for security, ecosystem health, and sustainable reuse. By doing so, we can move toward practices of reuse that are more intentional, balanced, and resilient—where developers not only benefit from shared code, but also understand the limits and responsibilities that come with dependency choices.